# High-Entanglement Capabilities for Variational Quantum Algorithms: The Poisson Equation Case

### v1

@@
The discretized Poisson equation matrix (DPEM) is vital to the field of computational fluid dynamics (CFD), and an algorithm that solves it on a quantum computer could potentially grant exponential space and time complexity speedups. However, the DPEM in 1D has been shown to have trouble being represented as a decomposition in the Pauli basis. Additionally, traditional ansatz for Variational Quantum Algorithms (VQAs) that are used to heuristically solve linear systems (such as the DPEM) have large numbers of parameters, making them harder to train. This research attempts to resolve these problems by utilizing the IonQ Aria quantum computer capabilities that boast all-to-all connectivity of qubits [1]. We propose a decomposition of the DPEM that is based on 2- or 3-qubit entanglement gates and is shown to have $O(1)$ terms with respect to system size, with one term having an $O(n^{2})$ circuit depth and the rest having only an $O(1)$ circuit depth (where $n$ is the number of qubits defining the system size). To test these new improvements, we ran numerical simulations to examine how well the VQAs performed with varying system sizes, showing that the new setup offers an $O(n)$ scaling of the number of iterations required for convergence, providing an exponential speedup over their classical computing counterparts. This project shows that the future of computational fluid dynamics may involve quantum computers to provide significant time and space complexity speedups.
@@ 1 Introduction
Additionally, in typical analysis of solving the DPEM with VQLS, the matrix itself is decomposed into a linear combination of sub-matrices, which has a number of terms that scales either exponentially [16] or polynomially [17] with respect to system size. This means that as the size of the matrix gets larger, the number of quantum circuits required to evaluate the cost function grows exponentially or polynomially. In this paper, we propose the usage of a high-entanglement decomposition (HED), which provides a constant scaling of sub-matrices with respect to system size, containing only four terms in the final decomposition [^1] The decomposition is inspired by the HED proposed in [16]. However, their decomposition is slightly incorrect and not optimal.. Additionally, three of the terms have an $O(1)$ circuit depth, and one term has an $O(n^{2})$ circuit depth, meaning that the HED does not require the evaluation of quantum circuits with circuit depth longer than the maximum on IonQ quantum computers [18].
@@ 2.1 Algorithm Overview
The VQLS uses a heuristic approach to solving the linear system $A\mathbf{x}=\mathbf{b}$. It begins by assuming that the matrix $A$ can be decomposed in some finite number of sub-matrices that can be represented as a combination and tensor product of quantum logic gates. Additionally, we assume that the vector $\mathbf{b}$ can be decomposed into a short-depth quantum circuit denoted as $U$, comprised of some sub-unitaries $U_{l}$, such that a quantum state $\left|b\right\rangle\propto\mathbf{b}$ can be prepared in constant circuit depth.
@@
We initially prepare some guess of the solution as a quantum state, which we call $\left|\psi(\boldsymbol{\theta})\right\rangle$, parameterized by a vector of rotations $\boldsymbol{\theta}$. This guess can be prepared using a parameterized quantum circuit known as a variational ansatz, in which the parameters $\boldsymbol{\theta}$ will determine certain rotation values of the circuit. The ansatz we use in this paper is a combination of parameterized rotation layers (using $R_{y}$ gates) and entanglement layers (using $CZ$ gates). An example of our ansatz for an $n=6$ system is shown in Figure 3.
@@ 2.2 The Poisson Equation
The most difficult part about the VQLS algorithm is the decomposition of the desired matrix into sub-matrices that can be represented on a quantum computer. For the DPEM, there are 2 decompositions we will discuss in this paper. The first utilizes the Pauli basis of quantum logic gates, which is easy to understand and has short circuit lengths when applied to the actual algorithm. The second decomposition is more complicated and requires quantum circuits that entangle almost all qubits with each other. However, this decomposition will be favorable in later use because its decomposition has an $O(1)$ scaling of sub-matrices and a generally well-behaved scaling of circuit depth, which is $O(n^{2})$. With currently available hardware (based on current circuit depth limitations of the IonQ Aria machine [20]), one could already simulate a one-million particle mesh with only twenty qubits using the HED method.
@@ 2.2.2 High Entanglement Decomposition
A second decomposition of the Poisson equation exists, although it requires more complicated procedures for its generation, and it requires that the quantum computer for which it runs on can complete two or three qubit entanglement gates between almost all qubits. IonQ’s latest quantum computers indeed are fully capable of such operations [20], so their quantum computers are a prime candidate for testing this decomposition of the matrix on a real quantum computer. The HED shortens the number of expectation values required to be evaluated per iteration, has a constant number of sub-matrices, and has reasonable circuit depth for each quantum circuit required to evaluate the cost function. This decomposition takes much less computational resources overall compared to the Pauli decomposition, so we use it in all our numerical simulations.
@@
which clearly has an $O(1)$ circuit depth. Figure 11(c)(a) shows the quantum circuit for $L_{1}$ for the $n=6$ case.
@@
Figure 11(b) depicts the quantum circuits representing $C_{1},C_{2},\ldots,C_{5}$ as an example for the $n=6$ case.
@@
An example of the implementation of this matrix as a quantum circuit is shown for the $n=6$ case in Figure 11(c)(b).
@@ 2.3 The Cost Function
2.3.1 Global Entangling Ansatz
In order to generate our $\left|\psi(\boldsymbol{\theta})\right\rangle$ state, we use a Global-Entangling Ansatz (GEA), which utilizes the capabilities of the IonQ hardware. The quantum computers available through IonQ’s quantum computing cloud services boast all-to-all connectivity between qubits, allowing one to easily encode two-qubit gates between any two qubits [1]. We define the GEA with multiple layers of parameterized rotation and entanglement gates, where the parameters are defined by $\boldsymbol{\theta}$. Additionally, we pre-conditioned the ansatz with the unitary that prepares the $\left|b\right\rangle$ state, denoted by $U_{b}$. We decided to see whether or not this type of entangling scheme would benefit the performance of the cost functions, and as shown through our numerical results, we found that it offered great improvements with respect to number of qubits and number of iterations when compared to the LEA (or HEA) proposed in [14]. There are certain factors going into how many layers (and thus the number of parameters) the ansatz should have, but simply put, if the parameter space is too small, the solution might not exist in it at all, but if it is too large, it might take far too long to find the solution [13]. The GEA used in this paper is defined with three repeated layers of rotation gates ($R_{y}$) and entangling gates ($CZ$), with a unique rotation parameter attributed to each $R_{y}$ gate. Thus, the number of parameters per variational ansatz is $3n$ where $n$ is the number of qubits defining the system size.
Figure 4: Parameters versus qubits depending on the ansatz chosen.
@@ 2.3.3 Locally and Globally Defined Cost
so our cost function $C_{G}$ may be defined as the expectation value of the following, given that $\Phi=A\left|\psi(\boldsymbol{\theta})\right\rangle$:
@@ 4.2 Statement of Funding
(b) $C_{1}$, $C_{2}$, …, and $C_{5}$ from left to right. Finally, (d) depicts the overall implementation of the $L_{2}$ matrix for $n=6$
(c) From left to right, the quantum circuits for $L_{1}$ and $L_{3}$
@@ Appendix B Remaining Numerical Simulations
(d) Remaining runs of $q_{\Delta}=0.01$ that were not shown in the figures above.
(e) Results of numerical simulations with $q_{\Delta}=0.1$. Here we see that the simulation times out after running simulations of system sizes larger than $n=4$, whereas the GEA does not have the same timeout problems. The simulations timed-out after iterations took longer than 12 hours to run. Simulations were run until they reached the precision metric defined in [14].

### v2

@@
The discretized Poisson equation matrix (DPEM) is vital to the field of computational fluid dynamics (CFD), and an algorithm that solves it on a quantum computer could potentially grant exponential space and time complexity speedups. However, the DPEM in 1D has been shown to have trouble being represented as a decomposition in the Pauli basis. Additionally, traditional ansatz for Variational Quantum Algorithms (VQAs) that are used to heuristically solve linear systems (such as the DPEM) have large numbers of parameters, making them harder to train. This research attempts to resolve these problems by utilizing the IonQ Aria quantum computer capabilities that boast all-to-all connectivity of qubits. We propose a decomposition of the DPEM that is based on 2- or 3-qubit entanglement gates and is shown to have $O(1)$ terms with respect to system size, with one term having an $O(n^{2})$ circuit depth and the rest having only an $O(1)$ circuit depth (where $n$ is the number of qubits defining the system size). To test these new improvements, we ran numerical simulations to examine how well the VQAs performed with varying system sizes, showing that the new setup offers an $O(n)$ scaling of the number of iterations required for convergence, providing an exponential speedup over their classical computing counterparts. This project shows that the future of computational fluid dynamics may involve quantum computers to provide significant time and space complexity speedups.
@@ 1 Introduction
Additionally, in typical analysis of solving the DPEM with VQLS, the matrix itself is decomposed into a linear combination of sub-matrices, which has a number of terms that scales either exponentially [16] or polynomially [17] with respect to system size. This means that as the size of the matrix gets larger, the number of quantum circuits required to evaluate the cost function grows exponentially or polynomially. In this paper, we propose the usage of a high-entanglement decomposition (HED), which provides a constant scaling of sub-matrices with respect to system size, containing only four terms in the final decomposition. The decomposition is inspired by the HED proposed in [16]. However, their decomposition is slightly incorrect and not optimal. Additionally, three of the terms have an $O(1)$ circuit depth, and one term has an $O(n^{2})$ circuit depth, meaning that the HED does not require the evaluation of quantum circuits with circuit depth longer than the maximum on IonQ quantum computers [18].
@@ 2.1 Algorithm Overview
The VQLS uses a heuristic approach to solving the linear system $A\mathbf{x}=\mathbf{b}$. The schematic of the algorithm can be seen in Figure 1. It begins by assuming that the matrix $A$ can be decomposed in some finite number of sub-matrices that can be represented as a combination and tensor product of quantum logic gates. Additionally, we assume that the vector $\mathbf{b}$ can be decomposed into a short-depth quantum circuit denoted as $U$, comprised of some sub-unitaries $U_{l}$, such that a quantum state $\left|b\right\rangle\propto\mathbf{b}$ can be prepared in constant circuit depth.
@@
We initially prepare some guess of the solution as a quantum state, which we call $\left|\psi(\boldsymbol{\theta})\right\rangle$, parameterized by a vector of rotations $\boldsymbol{\theta}$. This guess can be prepared using a parameterized quantum circuit known as a variational ansatz, in which the parameters $\boldsymbol{\theta}$ will determine certain rotation values of the circuit. The ansatz we use in this paper is a combination of parameterized rotation layers (using $R_{y}$ gates) and entanglement layers (using $CZ$ gates).
@@ 2.2 The Poisson Equation
The most difficult part about the VQLS algorithm is the decomposition of the desired matrix into sub-matrices that can be represented on a quantum computer. For the DPEM, there are 2 decompositions we will discuss in this paper. The first utilizes the Pauli basis of quantum logic gates, which is easy to understand and has short circuit lengths when applied to the actual algorithm. The second decomposition is more complicated and requires quantum circuits that entangle almost all qubits with each other. However, this decomposition will be favorable in later use because its decomposition has an $O(1)$ scaling of sub-matrices and a generally well-behaved scaling of circuit depth, which is $O(n^{2})$. With currently available hardware (based on current circuit depth limitations of the IonQ Aria machine [20]), one could already simulate a one-million particle mesh with only twenty qubits using the HED method, as shown in Figure 3(a).
@@ 2.2.2 High Entanglement Decomposition
A second decomposition of the Poisson equation exists, although it requires more complicated procedures for its generation, and it requires that the quantum computer for which it runs on can complete two or three qubit entanglement gates between almost all qubits. IonQ’s latest quantum computers indeed are fully capable of such operations [20], so their quantum computers are a prime candidate for testing this decomposition of the matrix on a real quantum computer. The HED shortens the number of expectation values required to be evaluated per iteration, has a constant number of sub-matrices, and has reasonable circuit depth for each quantum circuit required to evaluate the cost function, as will be shown in the following derivation and can be seen in Figure 3(a). This decomposition takes much less computational resources overall compared to the Pauli decomposition, so we use it in all our numerical simulations.
@@
which clearly has an $O(1)$ circuit depth. A.1(a) shows the quantum circuit for $L_{1}$ for the $n=6$ case.
@@
A.0(b) depicts the quantum circuits representing $C_{1},C_{2},\ldots,C_{5}$ as an example for the $n=6$ case.
@@
An example of the implementation of this matrix as a quantum circuit is shown for the $n=6$ case in Appendix A.1.
@@ 2.3 The Cost Function
2.3.1 Global Entangling Ansatz
In order to generate our $\left|\psi(\boldsymbol{\theta})\right\rangle$ state, we use a Global-Entangling Ansatz (GEA), which utilizes the capabilities of the IonQ hardware. The quantum computers available through IonQ’s quantum computing cloud services boast all-to-all connectivity between qubits, allowing one to easily encode two-qubit gates between any two qubits [15]. We define the GEA with multiple layers of parameterized rotation and entanglement gates, where the parameters are defined by $\boldsymbol{\theta}$. Additionally, we pre-conditioned the ansatz with the unitary that prepares the $\left|b\right\rangle$ state, denoted by $U_{b}$. We decided to see whether or not this type of entangling scheme would benefit the performance of the cost functions, and as shown through our numerical results, we found that it offered great improvements with respect to number of qubits and number of iterations when compared to the LEA (or HEA) proposed in [13]. There are certain factors going into how many layers (and thus the number of parameters) the ansatz should have, but simply put, if the parameter space is too small, the solution might not exist in it at all, but if it is too large, it might take far too long to find the solution [12]. As can be seen in Figure 3, the GEA has far fewer parameters than the HEA, which will make it easier to train. The high-entanglement nature of the GEA preserves its expressibility, thus making the solution to each system still well-accessible to the ansatz. The GEA used in this paper is defined with three repeated layers of rotation gates ($R_{y}$) and entangling gates ($CZ$), with a unique rotation parameter attributed to each $R_{y}$ gate. Thus, the number of parameters per variational ansatz is $3n$ where $n$ is the number of qubits defining the system size. An example of the GEA for $n=6$ can be found in Figure 4.
Figure 3: Parameters versus qubits depending on the ansatz chosen.
@@ 2.3.3 Locally and Globally Defined Cost
so our cost function $C_{G}$ may be defined as the expectation value of the projection Hamiltonian, given that $\Phi=A\left|\psi(\boldsymbol{\theta})\right\rangle$:
@@ 4.2 Statement of Funding
(b) $C_{1}$, $C_{2}$, …, and $C_{5}$ from left to right, which will be used to construct the circuit for the $L_{2}$ matrix.
A. 0: The final circuit that represents the application of the $L_{2}$ matrix to the ansatz.
A. 1: Quantum circuit representing $L_{1}$
A. 2: Quantum circuit representing $L_{3}$
@@ Appendix B Remaining Numerical Simulations
(a) Remaining runs of $q_{\Delta}=0.01$ that were not shown in the figures above.
(b) Results of numerical simulations with $q_{\Delta}=0.1$. Here we see that the simulation times out after running simulations of system sizes larger than $n=4$, whereas the GEA does not have the same timeout problems. The simulations timed-out after iterations took longer than 12 hours to run. Simulations were run until they reached the precision metric defined in [13].